Can someone explain how to do this in RQ5. (Exercise 1.49) Studies on treating Alzheimers disease. Alzheimers disease (AD) is a progressive disease of the brain. Much research has been conducted on how to treat AD The journal eCAM (November 2006) published an article that critiqued the quality of the methodology used in studies on AD treatment. For each in a sample of 13 studies, the quality of the methodology was measured using the Wong Scale, with scores ranging from 9 (low quality) to 27 (high quality). The data are shown in the table below. Estimate the mean quality, A, of all studies on the treatment of Alzheimers disease with a 99% confidence interval. Interpret the result. (1 pt) 22 21 18 19 20 15 19 20 15 20 17 20 21 Note: You can calculate the sample mean and the sample standard deviation by hand or using the following R code. (The Lab session note 1 will have the code to calculate the confidenoe interval.) Data = c(22,21,18,19,20,15,19,20,15,20,17,20,21) mean (Data) sd (Data)
Answer #1

> Data = c(22,21,18,19,20,15,19,20,15,20,17,20,21)
> m = mean(Data)
> m
[1] 19
> s = sd(Data)
> s
[1] 2.198484
> n = length(Data)
> n
[1] 13
> lc = m - qt(0.995,n-1) * s/sqrt(n)
> lc
[1] 17.13749
> uc = m + qt(0.995,n-1) * s/sqrt(n)
> uc
[1] 20.86251

The point estimate for mean is = 19

The 99% confidence interval is

(17.1375 , 20.8625)